# Making Offline RL Online: Collaborative World Models for Offline Visual Reinforcement Learning

### v3

@@ A.3 Behavior Learning
Appendix B Additional Quantitative and Qualitative Results
B.1 Visualizations on Policy Evaluation
We evaluate the trained agent of different models on the Meta-World and DMC tasks and select the first 45 frames for comparison. Figure 8 and Figure 9 present showcases of performing the learned policies of different models on DMC and Meta-World respectively.
(a) Policy evaluation on the DMC Walker Downhill task
(b) Policy evaluation on the DMC Walker Uphill task
(c) Policy evaluation on the DMC Walker Nofoot task
Figure 9: Policy evaluation on the Meta-World Button Topdown task. The performance of model-free method CURL is poor and it cannot complete the task (green box). CoWorld achieves better performance and completes the task in fewer steps (red box) than Offline DV2 (blue box).
B.2 Quantitative Results on DMC Meidum-Expert Dataset
Similar to the data collection strategy of medium-replay dataset, we build offline datasets of medium-expert quality using a DreamerV2 agent.
The medium-expert dataset comprises all the samples in the replay buffer during the training process until the policy attains expert-level performance, defined as achieving the maximum score that the DreamerV2 agent can achieve. As shown in Table 4, CoWorld outperforms other baselines on the DMC medium-expert dataset in most tasks.
Offline DV2 $450\pm 24$ $141\pm 1$ $214\pm 8$ $248\pm 9$ 3±$\pm 0$ $48\pm 3$ 184
DrQ+BC 808$\pm$47 762$\pm$61 808$\pm$45 862$\pm$13 454$\pm$12 730$\pm$17 737
LOMPO 548$\pm$245 449$\pm$117 688$\pm$97 174$\pm$29 19$\pm$10 113$\pm$35 332
Finetune 784$\pm$46 671$\pm$65 851$\pm$91 858$\pm$9 428$\pm$49 833$\pm$7 738
CoWorld 848$\pm$9 774$\pm$29 919$\pm$7 871$\pm$13 475$\pm$16 844$\pm$1 789
B.3 Quantitative Results on Meta-World
Figure 10(a) compares the performance of different models on Meta-World. DV2 Finetune demonstrates better performance in the initial training phase, thanks to its direct access to the source environment. Instead, CoWorld introduces auxiliary source value guidance to assist the training of the target agent. In the final phase of training, the source value guidance is more effective, and then CoWorld outperforms DV2 Finetune. Furthermore, Figure 10(b) presents the ablation studies of CoWorld conducted on Meta-World, highlighting the effectiveness and necessity of each training stage.
(a) Comparison with other baselines
(b) Ablation studies of CoWorld
B.4 Effect of Latent Space Alignment
(a) Latent space before alignment
(b) Latent space after alignment
B.5 Training Efficiency
As presented in Table 5, we evaluate the training/inference time on the Meta-World benchmark (Handle Press $\rightarrow$ Button Topdown) using a single RTX 3090 GPU.
Table 5: Time complexity on the Meta-World HP $\rightarrow$ BT task.
Model Number of Training Iterations Training Time Inference Time Per Episode
Offline DV2 300k 2054 min 2.95 sec
DrQ+BC 300k 200 min 2.28 sec
CQL 300k 405 min 1.88 sec
CURL 300k 434 min 2.99 sec
LOMPO 100k 1626 min 4.98 sec
DV2 Finetune 460k 1933 min 6.63 sec
DV2 Finetune+EWC 460k 1533 min 5.58 sec
CoWorld 460k 3346 min 4.47 sec
Appendix C Details of Multi-Source CoWorld
The key idea of adaptive domain selection of multi-source CoWorld is to allocate a set of one-hot weights $\omega_{t}^{i=1:N}$ to candidate source domains by calculating their KL divergence in the latent state space to the target domain, where $i\in[1,N]$ is the index of each source domain. The adaptive domain selection procedure includes the following steps:
  1. 1.
    World models pretraining. We pretrain a world model for each source domain and target domain individually.
  2. 2.
    Domain distance measurement. At each training step in the target domain, we measure the KL divergence between the latent states of the target domain, produced by $e_{\phi}(o_{t}^{(T)})$, and corresponding states in each source domain, produced by $e_{\phi^{\prime}_{i}}(o_{t}^{(T)})$. Here, $e^{(T)}_{\phi}$ is the encoder of the target world model and $e_{\phi^{\prime}_{i}}$ is the encoder of the world model for the source domain $i$.
  3. 3.
    Auxiliary domain identification. We dynamically identify the closest source domain with the smallest KL divergence. We set $\omega_{t}^{i=1:N}$ as a one-hot vector, where $\omega_{t}^{i}=1$ indicates the selected auxiliary domain.
  4. 4.
    Rest of training. With the one-hot weights, we continue the rest of the proposed online-to-offline training approach. For example, during representation learning, we adaptively align the target state space to the selected online simulator by rewriting the domain alignment loss term in Eq. (3) as
$\mathcal{L}_{\text{M-S}}=\beta_{2}\sum_{i=1}^{N}\omega_{i}\mathrm{KL}\left[\texttt{sg}(g(e_{\phi^{\prime}}(o_{t}^{(T)})))\ \|\ g(e_{\phi}(o_{t}^{(T)}))\right].$ (12)
To evaluate the effectiveness of the multi-source adaptive selection algorithm, we conducted experiments on Meta-World and RoboDesk Benchmark. For each target task, two source tasks are utilized, including the CoWorld best-performing task and the CoWorld worst-performing task. Additionally, the sub-optimal source task is added for some target tasks.
As shown in Table 6, multi-source CoWorld can adaptively select the best source task for most multi-source problems to ensure adequate knowledge transfer. The performance of multi-source CoWorld is reported in Table 2. CoWorld can flexibly adapt to the transfer learning scenarios with multiple source domains, achieving comparable results to the model that exclusively uses our manually designated auxiliary simulator as the source domain (best source). This study significantly improves the applicability of CoWorld in broader scenarios.
Table 6: Multi-source CoWorld selection results on Meta-World and RoboDesk.
Target Task Selected Source Task
Door Close Drawer Close
Button Press Handle Press
Window Close Button Topdown
Handle Press Button Press
Button Topdown Handle Press
Drawer Close Door Close
RoboDesk Push Button Button Press
RoboDesk Open Slide Window Close
RoboDesk Drawer Open Drawer Close
RoboDesk Upright Block off Table Handle Press
Appendix D Compared Methods
  • Offline DV2 (Lu et al., 2023): A model-based RL method that modifies DreamerV2 (Hafner et al., 2021) to offline setting, and adds a reward penalty corresponding to the mean disagreement of the dynamics ensemble.
  • DrQ+BC (Lu et al., 2023): It modifies the policy loss term in DrQ-v2 (Yarats et al., 2021) to match the loss given in (Fujimoto & Gu, 2021).
  • CQL (Lu et al., 2023): It is a framework for offline RL that learns a Q-function that guarantees a lower bound for the expected policy value than the actual policy value. We add the CQL regularizers to the Q-function update of DrQ-v2 (Kumar et al., 2020).
  • CURL (Laskin et al., 2020): It is a model-free RL approach that extracts high-level features from raw pixels utilizing contrastive learning.
  • LOMPO (Rafailov et al., 2021): An offline model-based RL algorithm that handles high-dimensional observations with latent dynamics models and uncertainty quantification.
  • LOMPO Finetune (Rafailov et al., 2021): It pretrains a LOMPO agent with source domain data and subsequently finetunes the pretrained agent in the offline target domain.
  • DV2 Finetune: It pretrains a DreamerV2 agent in the online source domain and subsequently finetunes the pretrained agent in the offline target domain. Notably, Meta-World $\rightarrow$ RoboDesk tasks’ action space is inconsistent, and we can’t finetune directly. Instead, we use the maximum action space of both environments as the shared policy output dimension. For Meta-World and Meta-World $\rightarrow$ RoboDesk transfer tasks, pretrain the agent 160k steps, and finetune it 300k steps. For DMC transfer tasks, pretrain the agent 600k steps, and finetune it 600k steps.
  • DV2 Finetune+EWC: It modifies the DV2 Finetune method with EWC to regularize the model for retaining knowledge from the online source domain. The steps of pretraining and finetuning are consistent with DV2 Finetune.
Appendix E Implementation Details
Meta-World.
RoboDesk.
In our study, we select Meta-World as the source domain and RoboDesk as the target domain. Notably, there exists a significant domain gap between these two environments. The visual observations, physical dynamics and action spaces of two environments are different. First, Meta-World adopts a side viewpoint, while RoboDesk utilizes a top viewpoint. Further, the action space of Meta-World is 4 dimensional, while RoboDesk comprises a 5-dimensional action space. Considering these differences, the Meta-World $\rightarrow$ RoboDesk presents a challenging task for transfer learning.
DeepMind Control.
Source agents are trained with standard DMC environments, target agents are trained in modified DMC environments. In this modified environment, Walker Uphill and Cheetah Uphill represent the task in which the plane is a $15^{\circ}$ uphill slope. Walker Downhill and Cheetah Downhill represents the task in which the plane is a $15^{\circ}$ downhill slope. We evaluate our model with baselines in six tasks with different source domains and target domains.
Appendix F Assumptions of the Similarity between the Source and Target Domains
We assume that there exist notable distinctions between the source and target domains (see Table 1). This assumption can be softened by our proposed approaches of state and reward alignment. Through these approaches, we aim to mitigate domain discrepancies between distinct source and target MDPs. Empirical evidence supporting our methods is presented in Section 4.3, where our proposed approach demonstrates robust performance in the Meta-World $\rightarrow$ RoboDesk transfer RL setup.
Our experiments reveal that the CoWorld method exhibits a notable tolerance to inter-domain differences in visual observation, physical dynamics, reward definition, or even the action space of the robots. This characteristic makes it more convenient to choose an auxiliary simulator based on the type of robot. For example:
  • When the target domain involves a robotic arm (e.g., RoboDesk), an existing robotic arm simulation environment (e.g., Meta-World as used in our paper) can be leveraged as the source domain.
  • For target domains related to autonomous driving vehicles, simulation environments like CARLA can be selected.
Appendix G Hyperparameters

### v4

@@ A.3 Behavior Learning
A.4 Hyperparameters
@@
Appendix B Additional Quantitative and Qualitative Results
B.1 Visualizations on Policy Evaluation
We evaluate the trained agent of different models on the Meta-World and DMC tasks and select the first $45$ frames for comparison. Figure 8 and Figure 9 present examples of performing the learned policies of different models on DMC and Meta-World respectively.
(a) Policy evaluation on the DMC Walker Downhill task
(b) Policy evaluation on the DMC Walker Uphill task
(c) Policy evaluation on the DMC Walker Nofoot task
Figure 9: Policy evaluation on the Meta-World Button Topdown task. The model-free method CURL cannot complete the task (green box). CoWorld achieves better performance and finishes the task in fewer steps (red box) than Offline DV2 (blue box).
B.2 Quantitative Results on DMC Medium-Expert Dataset
Similar to the data collection strategy of the medium-replay dataset, we build offline datasets with medium-expert quality using a DreamerV2 agent. The medium-expert dataset comprises all the samples in the replay buffer during the training process until the policy attains expert-level performance, defined as achieving the maximum score that the DreamerV2 agent can achieve. As shown in Table 9, CoWorld outperforms other baselines on the DMC medium-expert dataset in most tasks.
Offline DV2 $450\pm 24$ $141\pm 1$ $214\pm 8$ $248\pm 9$ 3±$\pm 0$ $48\pm 3$ 184
DrQ+BC 808$\pm$47 762$\pm$61 808$\pm$45 862$\pm$13 454$\pm$12 730$\pm$17 737
LOMPO 548$\pm$245 449$\pm$117 688$\pm$97 174$\pm$29 19$\pm$10 113$\pm$35 332
Finetune 784$\pm$46 671$\pm$65 851$\pm$91 858$\pm$9 428$\pm$49 833$\pm$7 738
CoWorld 848$\pm$9 774$\pm$29 919$\pm$7 871$\pm$13 475$\pm$16 844$\pm$1 789
B.3 Quantitative Results on Meta-World
Figure 10(a) compares the performance of different models on Meta-World. DV2 Finetune demonstrates better performance in the initial training phase, thanks to its direct access to the source environment. Instead, CoWorld introduces auxiliary source value guidance to assist the training of the target agent. In the final phase of training, the source value guidance is more effective, and then CoWorld outperforms DV2 Finetune. Figure 10(b) presents the ablation studies of CoWorld conducted on Meta-World, highlighting the effectiveness and necessity of each training stage.
(a) Comparison with other baselines
(b) Ablation studies of CoWorld
B.4 Effect of Latent Space Alignment
(a) Latent space before alignment
(b) Latent space after alignment
B.5 Additional Results on the Realistic Sim2Real Setup
Due to the limitation in experimental resources, we are unable to conduct experiments with real robots. We make efforts to construct a more realistic sim2real setup. The experiment is conducted with the identical robot control task for both the source and target domains. We manually introduce two types of noise into the visual observation and action space of the target domain, trying to mimic the complex and noisy real-world scenes.
  • Visual noise: We modify the original DeepMind Control environment by replacing the static background with dynamic backgrounds of random real-world videos.
  • Action noise: We add Gaussian noises $n_{t}$ sampled from $\mathcal{N}(0,1)$ to every dimension of the action in Meta-World and RoboDesk, which originally ranges in (-1,1). This mimics scenarios where the offline dataset is collected using a low-cost (less inaccurate) real robot.
As shown in Table 10, we compare CoWorld with the finetuned DreamerV2 model on this new setup. We apply noise of three magnitudes, $w\sim\{0.1,1,5\}$, in the Meta-World Button Topdown task, leading to noisy actions of $a_{t}^{\text{real}}=a_{t}+w\cdot n_{t}$.
Table 10: Results with more significant domain gaps.
      Target Domain       CoWorld       DV2 Finetune
      DMC Walker Walk       544       457
      DMC Cheetah Run       296       220
      RoboDesk Push Green ($w=1$)       406       358
      Meta-World Button Topdown ($w=0.1$)       3752       2693
      Meta-World Button Topdown ($w=1$)       3567       3104
      Meta-World Button Topdown ($w=5$)       951       670
From the above results, it is evident that CoWorld consistently outperforms the naive finetuning method in this ‘sim2real’ setup. Importantly, we assess the model under more challenging setups, with more significant domain gaps, as illustrated in Table 1.
B.6 Comparison with Pre-trained Foundation Model R3M
R3M [32] is pretrained on Ego4D human video dataset and facilitates efficient learning of downstream robotic tasks. R3M is shown to be a competitive model, particularly in its ability to transfer representations across domains with diverse visual inputs. For model comparison, we leverage the pre-trained weights of R3M from the official repository to initialize the representation model. We then perform policy optimization based on it for downstream tasks in the Meta-World environment. We respectively employ expert data, sourced from the official repository, alongside our own data, which is collected from scratch with mixed data quality. And DV2 Finetune is also pretrained in a related task and finetuned on the offline dataset. As demonstrated in Table 11, our approach outperforms the R3M / DV2 fine-tuning model.
Table 11: Comparison of CoWorld with using a pre-trained foundation model, R3M.
CoWorld R3M (expert data) R3M (our data) DV2 Finetune
Button Press Topdown 3889 1609 311 3499
Drawer Close 4845 N/A 4616 4273
Handle Press 4570 N/A 1603 3702
It is important to note that:
  • Despite its generalizable representations, R3M is NOT specifically designed to solve the value overestimation problem, which is fundamental in offline RL. In contrast, our approach not only aligns state representations across domains but also effectively tackles the issue of value overestimation, and therefore achieves better performance.
  • The fine-tuning process of R3M necessitates expert demonstrations for high-quality imitation learning. However, its performance empirically deteriorates when applied to the offline dataset of the medium-replay data.
  • The pre-training process of R3M typically takes around $5$ days on a V100 GPU, while the entire training procedure of our approach takes only about $2$ days on a 3090 GPU.
B.7 Training Efficiency
As shown in Table 12, we evaluate the training/inference time on Meta-World (Handle Press $\rightarrow$ Button Topdown) using a single RTX 3090 GPU. Empirically, CoWorld achieves convergence ($90\%$ of the highest returns) in approximately $14$ hours; while it costs DV2 Finetune about $13$ hours. These results indicate that CoWorld requires a comparable training wall-clock time to DV2 Finetune, while consistently maintaining better performance in terms of returns after model convergence.
Table 12: Runtime comparisons evaluated on Meta-World (HP $\rightarrow$ BT).
Model # Training iterations Training time Inference time per episode
Offline DV2 300k 2054 min 2.95 sec
DrQ+BC 300k 200 min 2.28 sec
CQL 300k 405 min 1.88 sec
CURL 300k 434 min 2.99 sec
LOMPO 100k 1626 min 4.98 sec
DV2 Finetune 460k 1933 min 6.63 sec
DV2 Finetune+EWC 460k 1533 min 5.58 sec
CoWorld 460k 3346 min 4.47 sec
Appendix C Multi-Source CoWorld
The key idea of multi-source CoWorld is to allocate a set of one-hot weights $\omega_{t}^{i=1:M}$ to candidate source domains by calculating their KL divergence in the latent state space to the target domain, where $i\in[1,M]$ is the index of each source domain. This procedure includes the following steps:
  1. 1.
    World models pretraining: We pretrain a world model for each source domain and target domain individually.
  2. 2.
    Domain distance measurement: At each training step in the target domain, we measure the KL divergence between the latent states of the target domain, produced by $e_{\phi}(o_{t}^{(T)})$, and corresponding states in each source domain, produced by $e_{\phi^{\prime}_{i}}(o_{t}^{(T)})$. Here, $e^{(T)}_{\phi}$ is the encoder of the target world model, and $e_{\phi^{\prime}_{i}}$ is the encoder of the world model for the source domain $i$.
  3. 3.
    Auxiliary domain identification: We dynamically identify the closest source domain with the smallest KL divergence. We set $\omega_{t}^{i=1:M}$ as a one-hot vector, where $\omega_{t}^{i}=1$ indicates the selected auxiliary domain.
  4. 4.
    Rest of training: With the one-hot weights, we continue the rest of the proposed online-to-offline training approach. During representation learning, we adaptively align the target state space to the selected online simulator by rewriting the domain alignment loss term in Eq. (3) as
$\mathcal{L}_{\text{M-S}}=\beta_{2}\sum_{i=1}^{M}\omega_{i}\mathrm{KL}\left[\texttt{sg}(g(e_{\phi^{\prime}}(o_{t}^{(T)})))\ \|\ g(e_{\phi}(o_{t}^{(T)}))\right].$ (12)
To evaluate the effectiveness of the multi-source adaptive selection algorithm, we conducted experiments on Meta-World and RoboDesk Benchmark. For each target task, two source tasks are used, including the CoWorld best-performing task and the CoWorld worst-performing task. Additionally, the sub-optimal source task is added for some target tasks.
As shown in Table 13, multi-source CoWorld can adaptively select the best source task for most multi-source problems to ensure adequate knowledge transfer. The performance of multi-source CoWorld is reported in Table 2. CoWorld flexibly adapts to the transfer learning scenarios with multiple source domains, achieving comparable results to the model that exclusively uses our manually designated auxiliary simulator as the source domain (best-source). This study significantly improves the applicability of CoWorld in broader scenarios.
Table 13: The source domain automatically selected by Multi-Source CoWorld. MW represents Meta-World and RD stands for RoboDesk.
         Target domain          Selected source domain
         MW: Door Close          MW: Drawer Close
         MW: Button Press          MW: Handle Press
         MW: Window Close          MW: Button Topdown
         MW: Handle Press          MW: Button Press
         MW: Button Topdown          MW: Handle Press
         MW: Drawer Close          MW: Door Close
         RD: Push Button          MW: Button Press
         RD: Open Slide          MW: Window Close
         RD: Drawer Open          MW: Drawer Close
         RD: Upright Block off Table          MW: Handle Press
Appendix D Source and Target Domains
Meta-World.
RoboDesk.
We select Meta-World as the source domain and RoboDesk as the target domain. Notably, there exists a significant domain gap between these two environments. The visual observations, physical dynamics, and action spaces of the two environments are different. First, Meta-World adopts a side viewpoint, while RoboDesk uses a top viewpoint. Further, the action space of Meta-World is $4$ dimensional, while that in RoboDesk is $5$-dimensional. Considering these differences, the Meta-World $\rightarrow$ RoboDesk benchmark presents a challenging transfer learning problem.
DeepMind Control.
We train the source agents in standard DMC environments and train the target agents in modified DMC environments. Walker Uphill and Cheetah Uphill represent tasks in which the ground has a $15^{\circ}$ uphill slope. Walker Downhill and Cheetah Downhill represent the tasks in which the plane has a $15^{\circ}$ downhill slope. We evaluate the model in six tasks with different source domains and target domains.
We assume that there exist notable distinctions between the source and target domains (see Table 1). This assumption can be softened by our proposed approach that mitigates domain discrepancies between distinct source and target MDPs. Our experiments reveal that the CoWorld method exhibits a notable tolerance to inter-domain differences in visual observation, physical dynamics, reward definition, or even the action space of the robots. This characteristic makes it more convenient to choose an auxiliary simulator based on the type of robot. For example:
  • When the target domain involves a robotic arm (e.g., RoboDesk), an existing robotic arm simulation environment (e.g., Meta-World) can be leveraged as the source domain.
  • For target domains related to autonomous driving, simulators like CARLA can be used.
Appendix E Compared Methods
  • Offline DV2 [25]: A model-based RL method that modifies DreamerV2 [16] to offline setting, and adds a reward penalty corresponding to the mean disagreement of the dynamics ensemble.
  • DrQ+BC [25]: It modifies the policy loss term in DrQ-v2 [50] to match the loss given in [10].
  • CQL [25]: It is a framework for offline RL that learns a Q-function that guarantees a lower bound for the expected policy value than the actual policy value. We add the CQL regularizers to the Q-function update of DrQ-v2 [21].
  • CURL [22]: It is a model-free RL approach that extracts high-level features from raw pixels utilizing contrastive learning.
  • LOMPO [39]: An offline model-based RL algorithm that handles high-dimensional observations with latent dynamics models and uncertainty quantification.
  • LOMPO Finetune: It pretrains a LOMPO agent [39] with source domain data and subsequently finetunes the pretrained agent in the offline target domain.
  • DV2 Finetune: It pretrains a DreamerV2 agent [16] in the online source domain and subsequently finetunes the pretrained agent in the offline target domain. Notably, Meta-World $\rightarrow$ RoboDesk tasks’ action space is inconsistent, and we can’t finetune directly. Instead, we use the maximum action space of both environments as the shared policy output dimension. For Meta-World and Meta-World $\rightarrow$ RoboDesk transfer tasks, we pretrain the agent for $160$k steps and finetune it $300$k steps. For DMC transfer tasks, we pretrain the agent for $600$k steps and finetune it for $600$k steps.
  • DV2 Finetune+EWC: It modifies the DV2 Finetune method with EWC [19] to regularize the model for retaining knowledge from the online source domain. The steps of pretraining and finetuning are consistent with DV2 Finetune.
Appendix F Broader Impacts
In real-world scenarios of healthcare applications, Zhang et al. [62] employed offline RL algorithms to train policies using a large amount of historical dataset, determining the follow-up schedules and tacrolimus dosages in Kidney Transplantation and HIV. There are also corresponding simulators [17, 1] designed by medical domain experts, with parameters learned from real-world data.
Another practical use of the proposed setup is advertising bidding, where direct interactions with real online advertising systems for training are challenging. A recent solution involves constructing a simulated bidding environment based on historical bidding logs for interactive training, such as [30], and mitigating the inherent differences between the virtual advertising environment and real-world advertising systems. Therefore, in many real-world scenarios, it is possible to optimize the policies learned from offline datasets with simulators.